OpenStudy (anonymous):

The dimensions of Eric’s pool are 9 feet by 7 feet. He wants to build a walkway surrounding the pool so that the square footage of the walkway and pool is 143 square feet. How wide is one side of the walkway? Show Work.

OpenStudy (anonymous):

dear i feel it will be 144 square feet instead of 143 square feet

OpenStudy (anonymous):

length=9ft + {walkway on both sides}=9+2x width=7ft + {walkway on both sides}=7+2x 143=(9+2x)(7+2x) 0=4x^2+32x-80 0=(x-2)(x+10) x=-10,2 since we can't have a negative length for the walkway, it must be positive. hence our walkway is 2ft wide. area={length}*{width} ={9+2x}*{7+2x} ={13}*{11} =143

OpenStudy (anonymous):

what about square footage including pool???????